OpenStudy (anonymous):

using multiplication (show work) 1) -10 = k/15 2) -56/m = -8

OpenStudy (anonymous):

and where's the question mark?

OpenStudy (anonymous):

? <- there

OpenStudy (anonymous):

Its a shame you don't even know how to use it. Make your question clear and maybe you'll get some answers. Otherwise stop wasting people's time.

OpenStudy (radar):

-10=k/15 multiply both sides by 15 -150=k you see how it was solved by using multiplication. Do you want to try the next one? I will stand by and review your work.

OpenStudy (radar):

Here is the 2nd problem: -56 ------- = -8 m What would I multiiply both sides by to get m to one side or the other and no longer have it showing in the denominator. Have you learned about "cross multiplying" yet?

OpenStudy (anonymous):

no I haven't learned about cross multiplying. Multiply both sides by -8?

OpenStudy (radar):

Good guess but we don't want to multiply both sides by -8

OpenStudy (anonymous):

oh hahah whoops. would it be 7 then, because 56 divided by 8 is 7?

OpenStudy (radar):

We would still have m on the bottom (in the denominator) We need to multiply both sides by m! I know that seems strange but here is what we get when we multiply both sides by m. -56=-8m Do you follow that catfish312?

OpenStudy (anonymous):

I do follow :)

OpenStudy (radar):

Now we could divide both sides by -8, but they wanted you to use multiplication. So we are going to multiply both sides by -1/8 getting 7=m or m=7. Do you see how multiplying by 1/8 is the same as dividing by 8?

OpenStudy (radar):

Those were the answers to those two problems, and we abided by the rules....... we multiplied .lol

OpenStudy (anonymous):

Thanks radar! :)

OpenStudy (radar):

You're welcome and good luck with these, this practice is getting you ready for bigger things to come.

OpenStudy (anonymous):

1) -10 = k/15 --> k = 15 *(-10) = - 150 2) -56/m = -8 --> m = -56/-8 = 7